day shift summary for 06/16/99

AG (ag@fnal.gov)
Wed, 16 Jun 1999 16:37:18 -0400 (EDT)

Shift Summary for day shift, 06/16/99

On Shift: Arun ( AT ), Sasha ( AG )

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of E832 data (special runs not included): ~7.
Hours of special runs (laser scans, muon runs...): ~1.5

Accounting of lost time:

Adjustment of HV's on some CsI channels with shifted HCC thresholds.

========================================================================
Runs ended this shift:

Run 13623 - 7.5M events - BEAM_832
Run 13624 - PED
Run 13625 - 0.9M - BEAM_832
Run 13626 - COSMIC
Run 13627 - running - BEAM_832

Tapes brought to FCC this shift: QKC010-12, QKK010-12, QKS010-12,
(including tapes for online split) QKB004, QKSR02, QKMU05, QKPE01

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

Goal(s) of shift: 1. E832 data-taking

2. Make ped run to check if reg peds (crate 15)
are still off by +5 counts.

3. If no beam, then make long laser scan.

Run 13623 finished normally. We made ped run and found all reg
peds back to normal. New e8323 run 13624 started, but we noticed
that HCC thresholds were way off, so CsI HV's had to be adjusted.
Cosmic run 13625 had been started for the HV's adjustment.
After the HV adjustment, database values were updated and new e832
run had been started.

RickF reduced the horizontal spot size of the beam to ~8 wires.

We checked on-line histos and found some bugs/problems.
People are notified. Also, many histograms in the folder
have to be updated.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13623 149 0828 everything is normal

163 0847 beam delayed to EBL due to frig permit,
est is 10 min.

169 0855 beam is back

193 0927 ktevana errors, ran kfrend

257 1053 tapes are full, run stops itself.
unload tapes, put new ones in.
drt_fresh_start before loading new tapes.

13624 1 1111 start ped run to check reg peds
now it looks that they are all back
to normal. the conclusion is that
something was wrong with the previuos
ped run.

4 1120 stop the run, update peds.

13625 1 1130 start another e832 run.

12 1141 SEM is 8.9e12; called MCR; they'll tune it

14 1145 SEM is back to normal: 4.6e12

26 1200 stop the run

13626 1 1210 start cosmic run to adjust HV on some
CsI channels

46 1300 stop the run. HV's have been adjusted and
constants are updated.

13627 1 1310 start another e832 run

1427 ktevana errors, ran kfrend.
SY energy conservation - no beam

61 1437 beam is back

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Page Arun (722 6855) or call Elizabeth (3399)
if there is a downtime of over 1.5 hrs.
Need to change some BA channels to another shaper-splitter module.

Take a laser scan if beam is off for an hour or so.